OpenStudy (anonymous):

What is the zero of (2-x)?

OpenStudy (anonymous):

2-x=0 x=2-0 x=2 Answer: 2.

OpenStudy (anonymous):

Thank you ;)) how about (3+x) ?

OpenStudy (anonymous):

3+x=0 x=0-3 x=-3 Answer: -3.